With a world population in 2003 of 6.314 billion and a projected population growth rate of 1.1% instead ( which is the united nation’s projection for the period 2003 to 2015)
a. Compute the doubling time
b. State the year in which the world’s population would be 12.628 billion
1
Expert's answer
2020-10-20T06:28:15-0400

a) Using the rule of 70 for getting the doubling time ;

="\frac{70}{Rate of growth }", where r is the rate of growth.


="\frac{70}{1.1}" =63.63 yrs


Hence it will have a doubling time of 64 years


b)

"\frac{12.628}{6.314}=2"


This shows that its the doubling time hence


=2003+64 yrs =2067


=2067